OpenStudy (anonymous):

g(x)=square root of x-5 .. fins the domain and sketch the graph..

OpenStudy (anonymous):

\[\sqrt{x-5}\] is the domain \[x>5\]

OpenStudy (nory):

Well, you can't find a square root of a negative number, so the values that make the expression under the radical negative are not in the domain.

OpenStudy (nory):

Actually, you want x>=5, because x can be 5. You can take the square root of 0.

OpenStudy (anonymous):

i see..

OpenStudy (anonymous):

how about sketching the graph???

OpenStudy (nory):

Do you know about finding the inverse function?

OpenStudy (anonymous):

cant remember :/

OpenStudy (nory):

This is a very bad description, but basically you take a function and switch the x and y values to get its inverse.

OpenStudy (anonymous):

is this a inverse function?

OpenStudy (nory):

So, for example, if I had a function that had (4,5) as a pair, I'd switch it, and the corresponding coordinates for the inverse function would be (5,4).

OpenStudy (anonymous):

ok ..

OpenStudy (nory):

This function is the inverse of (x-5)^2 = y. Can you plot some points on that graph?

OpenStudy (nory):

The little ^2 means "squared."

OpenStudy (nory):

Actually, something is wrong with my explanation. So just plot some points on the graph of the original function and see what you get.

OpenStudy (nory):

Doesn't it look a bit like a parabola? That's what I always think.

OpenStudy (anonymous):

\[\sqrt{x-5}\] is same as \[\left( x-5 \right)^2\] ?????????????

OpenStudy (nory):

No, it isn't. Sorry about that. Try plotting some points, like I suggested above.

OpenStudy (anonymous):

well, this is what i got if x=6, y=1 ...

OpenStudy (anonymous):

is it correct?

OpenStudy (anonymous):

do you know how \(\large y=\sqrt{x} \) looks like?

OpenStudy (anonymous):



OpenStudy (anonymous):

no...